OpenStudy (anonymous):

A right triangle has a hypotenuse of length 20 and a leg of length 12. What is the length of the other leg? If necessary, round your answer to two decimal places.

OpenStudy (anonymous):

16

OpenStudy (anonymous):

L^2 = 20^2 - 12^2 = 400-144 = 256 hence L = 16